Find both the x-intercept and the y-intercept of the line given by the equation
2x+6y-14=0

Answer by Alan3354(69443)

Find both the x-intercept and the y-intercept of the line given by the equation
2x+6y-14=0
---------------
x+3y=7 (reduced)
To find the y-intercept, set x=0.
0+3y=7
y = 7/3 --> (0,7/3)
----------------
The line intercepts the x-axis (x-intercept) when y = 0
x+3*0 = 7
x = 7 --> (7,0)